OpenStudy (amonoconnor):

A question pertaining to Indefinite Integrals, and Anti-Deriviatives: My problem: "integral[((9r^2)(dr))/(sqrt.(1-r^3))] , u = (1- r^3)" **...I know that's ugly... I'm using Safari on an iPad, which goofs up some OS features like inserting pictures, and the equations toolbar thing. Sorry! :( So, in school, we are evaluating these using "U Substitution", and I am wondering if: 1. The denominator can be written as to the negative half power, and come up. 2. ...what do I do with the "r^2" in "9r^2", if u = (1- r^3)? Any and all help is greatly appreciated! :)

OpenStudy (amonoconnor):

I have a fair understanding of these problems (started them today in Calc), but this one is stumping me. If someone could show the work of them working through the problem, step by step, that would be the most beneficial, and greatly appreciated! (I am a very visual learner, and seeing the process really helps me understand what I'm doing)

ganeshie8 (ganeshie8):

\(\large \int \frac{9r^2 dr}{\sqrt{1-r^3}}\)

ganeshie8 (ganeshie8):

are u able to see latex on ur ipad correctly ?

OpenStudy (amonoconnor):

Yes!!

ganeshie8 (ganeshie8):

Good :) U-substitution involves 3 steps

OpenStudy (amonoconnor):

...sorry, I just meant that what you posted was the right integral... What's latex?

ganeshie8 (ganeshie8):

nvm, i see ipad is processing latex correctly :)

ganeshie8 (ganeshie8):

step1 : make some substitution (in present example, \(u = 1-r^3\) )

ganeshie8 (ganeshie8):

we're done wid step1 already, right ?

OpenStudy (amonoconnor):

Yes, and I understand that much:)

ganeshie8 (ganeshie8):

step2 : find \(du\) in terms of \(dr\)

ganeshie8 (ganeshie8):

lets see what it means by working out \(du\) in present problem

OpenStudy (amonoconnor):

Is it: du = ( 0 - 3r^2) dr ?

ganeshie8 (ganeshie8):

Yes !

ganeshie8 (ganeshie8):

\(\large \int \frac{9r^2 dr}{\sqrt{1-r^3}}\) substitute \(\large u = 1-r^3\) that gives \(\large du = -3r^2 dr\) \(\large \implies r^2dr = \frac{-du}{3} \)

ganeshie8 (ganeshie8):

plug them in ur integrand

OpenStudy (amonoconnor):

...can the denominator be written as (1-r^3)^(-1/2) ?

OpenStudy (amonoconnor):

As if multiplied to 9r^2?

ganeshie8 (ganeshie8):

you can do that

ganeshie8 (ganeshie8):

\(\large \int \frac{9r^2 dr}{\sqrt{1-r^3}}\) substitute \(\large u = 1-r^3\) that gives \(\large du = -3r^2 dr\) \(\large \implies r^2dr = \frac{-du}{3}\) then the integrand becomes : \(\large \int \frac{9(\frac{-du}{3})}{\sqrt{u}}\) after some simplification : \(\large -3\int u^{\frac{-1}{2}}du\)

ganeshie8 (ganeshie8):

^^let me knw if smthng doesnt make sense

OpenStudy (amonoconnor):

*working it out on my paper... I'll let you know:)

ganeshie8 (ganeshie8):

okie :)

OpenStudy (amonoconnor):

...I get what you did, and I follow, but just for clarification : why did you not, when "isolating" dr, divide by -3r^2, and only by -3??

ganeshie8 (ganeshie8):

you could do that if u prefer

ganeshie8 (ganeshie8):

I didnt divide by entire \(-3r^2\), cuz i saw that the numerator has \(r^2dr\) term already so, i thought of replacing \(r^2 dr\) with something

OpenStudy (amonoconnor):

Ohhh... I see... You worked smart, not hard:)

ganeshie8 (ganeshie8):

hahah, next u knw what needs to be done ?

OpenStudy (amonoconnor):

Yep, u^(-1/2) = u^1/2 + C so, the final answer is "3u^1/2 + C" ?

OpenStudy (amonoconnor):

*plz be right..

ganeshie8 (ganeshie8):

nope

OpenStudy (amonoconnor):

. . . Rage. . .

ganeshie8 (ganeshie8):

\(\large \int x^n dx = \frac{x^{n+1}}{n+1} + c\)

OpenStudy (amonoconnor):

Oh! Do we plug 1-r^3 back in ?

OpenStudy (amonoconnor):

Oh, yeah, I forgot about that part... *more coffee

ganeshie8 (ganeshie8):

\(\large \int \frac{9r^2 dr}{\sqrt{1-r^3}}\) substitute \(\large u = 1-r^3\) that gives \(\large du = -3r^2 dr\) \(\large \implies r^2dr = \frac{-du}{3}\) then the integrand becomes : \(\large \int \frac{9\frac{-du}{3}}{\sqrt{u}}\) after some simplification : \(\large -3\int u^{\frac{-1}{2}}du \) this integral is easy to evaluate now, evaluating it gives : \(\large -3\frac{u^{\frac{-1}{2}+1}}{\frac{-1}{2}+1} + c \) \(\large -6 u^{\frac{1}{2}} + c \) \(\large -6 (1-r^3)^{\frac{1}{2}} + c \)

ganeshie8 (ganeshie8):

yes, you're right! we need to substitute back the "u" s

ganeshie8 (ganeshie8):

"u" is just a passing cloud, it was not present in original integral, we substituted it in the middle... and so we should replace it by r's in the end

OpenStudy (amonoconnor):

Jesus, you're amazing... :) thank you so much!! You're a lifesaver, once again! I have one more problem on my assignment for tonight, and it looks pretty psycho... Will you be on OS for a while? In case I need your brilliance to guide my understanding again? I want to try it myself, but just in case...